OpenStudy (anonymous):

Were finding the slope intercept form of the equation.. i need help on, -13/6=-1/6y + 1x if anyone understand how to do this please help meee;p thanks! :)

OpenStudy (ash2326):

We have \[\frac{-13}{6}=-\frac{1}{6}y+x\] A slope intercept form of equation is of the form \[y=mx+c\] m=slope of the line c=y-intercept of the line Here first step will be to multiply both sides by 6. @power98 Will you do that and post what do you get?

OpenStudy (anonymous):

1/6y=x+13/6 ? thats right isnt it?

jimthompson5910 (jim_thompson5910):

What ash2326 is saying is this -13/6=-1/6y + 1x 6(-13/6) = 6(-1/6y + 1x) 6(-13/6) = 6(-1/6y) + 6(1x) -13 = -y + 6x